Database last updated on Sun, 01 Feb 2026 16:10:52 +0000 by localhost

Instrument constant changes for the week before 2024015 (2 sec, 0 rows)

SL burned out? (AP results when SL is ON) (0 sec, 6 rows)

HG burned out? (AP results when HG is ON) (0 sec, 16 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
02620240153919.0

Non-critical warnings from b-files' comments for days between 2024015 and 2024015 (0 sec, 4 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
y:Command file y not found299B01524.026TORON#02610:53:42found1-1
hg:HOME key pressed499B01524.026TORON#02610:56:411-1
checkhg:Last HG had to adjust the micrometer by at least 1399B01524.026TORON#02612:34:161-1
pd:Going to menu199B01524.026TORON#02616:21:421-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhgrssl
015TORONTO#026menu,Must do HG now,Waiting until ZA = 170 now 124.7,Waiting until ZA = 170 now 126.5,Waiting until ZA = 170 now 128.3,Waiting until ZA = 170 now 130.1,Waiting until ZA = 170 now 131.9,Waiting until ZA = 170 now 133.7,Waiting until ZA = 170 now 135.4,Waiting until ZA = 170 now 137.2,Waiting until ZA = 170 now 138.9,Waiting until ZA = 170 now 140.6,Waiting until ZA = 170 now 142.2,Waiting until ZA = 170 now 143.8,Waiting until ZA = 170 now 145.4,Waiting until ZA = 170 now 146.9,Waiting until ZA = 170 now 148.4,Waiting until ZA = 170 now 149.8,Waiting until ZA = 170 now 151.1,Waiting until ZA = 170 now 152.4,Waiting until ZA = 170 now 153.5,Waiting until ZA = 170 now 154.5,Waiting until ZA = 170 now 155.3,Waiting until ZA = 170 now 156.1,Waiting until ZA = 170 now 156.6,Waiting until ZA = 170 now 157,Waiting until ZA = 170 now 157.2,Waiting until ZA = 95 now 93.6626056
(1 sec)

Dead time (0 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B01524.0266#0263130